dude:

Displacement, velocity, speed, acceleration part 2

dude:

1 attachment
dude:

We want v(t)=0 Finding v(t) \(v(t)=3t^2-18t+24\)

dude:

\(3t^2-18t+24=0\) \(3(t-2)(t-4)=0\) So we have at \(t=2,4\) Then to find acceleration, are 2 and 4 on the right track? ;b

AZ:

so set it equal to 0 and solve and then we'll evaluate what the acceleration is at those times when velocity = 0

AZ:

On the right track, champ!

dude:

<3 Acceleration: \(a(t)=6t-18\) When t is 2 and 4 \(a(2)=6(2)-18=-6\) \(a(4)=6(4)-18=6\)

surjithayer:

\[acceleration =\frac{ dv }{dt }=6t-18\]

surjithayer:

at t=2 ,it is retardation

dude:

XDD

AZ:

haha and then similarly for part B, just set a(t) = 0 and then determine the speed at those times

dude:

Now for speed when acceleration is 0 \(6t-18=0\) \(t=3\) Then to speed \(|v(t)|=|3(3)^2-18(3)+24|=3m/s\)

AZ:

splendid

dude:

For total distance \(|s(3)-s(2)|+|s(2)-s(1)|+|s(1)-s(0)|\)? Or is there another strat

surjithayer:

https://www.desmos.com/calculator/9vglnfohey find distance from 0 to 2 and 2 to 3,then add

AZ:

since v(t) = 0 at t = 2,4 Those are the time when it changes directions that means the two intervals we're going to have from t = 0 to t =3 is going to be 0 to 2 and 2 to 3

surjithayer:

from the graph we can conclude s=20+2=22

AZ:

Otherwise, the other strategy once you cover integrals is \[\text{distance travelled} = \int\limits_{a}^{b}~|v(t)|~dt\] where a and b are the time intervals

AZ:

and displacement would be \[\text{displacement} = \int\limits_{a}^{b}~v(t)~dt\]

dude:

ty ty

AZ:

yw yw I will add that when you're integrating, you have to separate the bounds whenever it changes direction similar to what we did so you would have two integrals to solve and it would give us the same answer \[\text{distance} = \int\limits_{0}^{2}~(3t^2−18t+24)~dt + \int\limits_{2}^{3}~(3t^2−18t+24)~dt\] \(\text{distance} = 20 + 2 = \boxed{22}\)